# The Binary Fraction of Stars in the Dwarf Galaxy Ursa Minor via Dark Energy Spectroscopic Instrument

### v1

@@ 4.1 Binary Fraction in UMi
Figure 7: We present the PPDs for the metal-rich (orange) and metal-poor (green) populations, as indicated in the legend. The top and bottom panels are based on the 1991A&A...248..485D and 2017ApJS..230...15M models, respectively. The blue curves show the results for the full sample (as in Figure 6).
4.2 Selection Effects
In Section 4.1, we report a marginal indication that the metal-rich stellar population of UMi may exhibit a higher binary fraction than the metal-poor population. In this section, we examine whether this metallicity dependency could arise from sample selection effects rather than reflecting an intrinsic difference. Specifically, we investigate the impact of non-uniform observing strategies across UMi, including variations in the number of epochs and the time baselines or separations between them. For example, we know that the spatial distributions of metal-rich and metal-poor stars differ, with more metal-rich stars in central regions. The tiling of UMi, however, is spatially varying (see Figure 1), which may result in different number of epochs and time baselines for different stellar populations and may affect the comparison between different populations.
@@ 4.2 Selection Effects
Figure 8: Resampling of the metal-rich and metal-poor populations to match their distributions in number of epochs (left column), distance to the dwarf center in projection (semi-major axis of elliptical isophotes in number density, middle column), and LOSV uncertainty (right column), respectively. Top: Original distributions for the metal-rich (orange) and metal-poor (green) samples, along with the matched distribution after resampling (dashed red). Middle: The PPDs of the binary fraction for the 1991A&A...248..485D model following resampling in both populations. Bottom: The corresponding results for the 2017ApJS..230...15M model. We perform the resampling procedure 100 times. The translucent histograms show the distribution of results from individual realizations, while the bold lines represent the combined PPDs from all trials.
@@
The middle (1991A&A...248..485D model) and bottom (2017ApJS..230...15M model) rows of Figure 8 show the PPDs of the binary fractions for the resampled metal-rich (orange) and metal-poor (green) populations. Note that the solid curve is the overall PPD from the 100 resampled realizations mentioned above, with the results for the 100 realizations shown in the background by curves with thin, lighter colored lines. Although the resampling slightly alters the measured binary fractions and increases the uncertainties due to the reduced sample size, the difference between the binary fractions of the metal-rich and metal-poor populations remains.
@@
Figure 9: Similar to Figure 8, we resample both the metal-rich and metal-poor subsamples to ensure they share the same distribution of time separations across all observation pairs. Top: The distribution of the time separation in units of day for individual pairs of observations. Middle: The PPDs derived from the 1991A&A...248..485D model after resampling. The orange histogram is the result of the metal-rich population, while the green one is the result of the metal-poor population. Bottom: The corresponding results from the 2017ApJS..230...15M model. Both models show consistent results, with the discrepancy between the two subsamples persisting despite the resampling.
@@
4.3 Systematics Associated with the Period Distribution
@@
In addition to observational factors, the inferred binary fraction is highly sensitive to the choice of model parameter distributions. As mentioned in Section 3.1, the orbital period distribution plays a major role in constraining the binary fraction, with a lower mean value leading to a substantial decrease in the inferred binary fraction.
@@
In Section 4.1, the binary fraction derived using the 2017ApJS..230...15M model is slightly higher than that obtained with the 1991A&A...248..485D model. The key distinction between these models lies in their treatment of orbital parameter distributions. While the 1991A&A...248..485D model assumes independent orbital parameter distributions, the 2017ApJS..230...15M model provides joint distributions that account for correlations between different orbital parameters. We present the period distributions in Figure 10 for different models. Compared to the 1991A&A...248..485D model, represented by the orange line, the 2017ApJS..230...15M model, shown by the blue line, predicts a bit smaller fraction of short-period binaries. This is the major reason that leads to an overall slight increase in the inferred binary fraction by the 2017ApJS..230...15M model, as demonstrated in our results above.
@@
This sensitivity is also demonstrated in 2018AJ....156..257S, which infers that the binary fraction decreases as the mean period value of the model, $\mu_{\log P}$, increases. 2018AJ....156..257S uses three different models, including 1992ApJ...396..178F (hereafter 1992ApJ...396..178F), 1991A&A...248..485D and 2011MNRAS.417.1702M, which differ only in the assumed mean period values of $\mu_{\log P}$ = 3.5, 4.8, and 5.8, respectively. The corresponding binary fractions derived from these models by 2018AJ....156..257S are 0.52, 0.78, and 0.96, illustrating the strong dependency of the inferred binary fraction on the choice of period distribution.
@@
To further investigate the impact of the orbital period distribution, we repeat our constraints adopting the mean period value from the 1992ApJ...396..178F model [^5] We have also explored the 2011MNRAS.417.1702M period distribution model, which leads to a binary fraction very close to one and might not be realistic, so we do not show the results based on this model., which is lower than the value of 1991A&A...248..485D, while keeping all other orbital parameter distributions identical to those in 1991A&A...248..485D. For clarity, we refer to this modified version as the “1992ApJ...396..178F” model for our discussions afterward. The green line in Figure 10 represents the orbital period distribution for the 1992ApJ...396..178F model, which shows a higher fraction of shorter-period binaries compared to the 1991A&A...248..485D model. As shown in the upper panel of Figure 11, the binary fraction derived from the 1992ApJ...396..178F model decreases by approximately 0.2 for the full sample compared with that from 1991A&A...248..485D. This difference arises primarily due to the shift towards shorter orbital periods, and thus a lower binary fraction is preferred for the same LOSV variability.
@@
The bottom panel of Figure 11 shows that the binary fractions for both the metal-rich and metal-poor subsamples decrease by approximately the same amount when applying the 1992ApJ...396..178F model with a lower mean period. Despite the decrease in the inferred binary fraction, the discrepancy between the metal-rich and metal-poor populations persists (see the cyan and purple curves). Here we emphasize that we are applying the same orbital period distribution model to both metal-rich and metal-poor populations. If the metal-rich and metal-poor populations follow different period distributions, such as a shorter mean period for the metal-rich population and a longer mean period for the metal-poor population, the observed discrepancy in binary fractions could be reduced or even eliminated. This effect can be seen by comparing the green curve of the metal-poor population and the cyan curve for the metal-rich population in the bottom panel of Figure 11. However, explaining the discrepancy in this way would require the metal-rich and metal-poor populations to have fundamentally different orbital period distributions, for which there is no strong theoretical support and which is therefore unlikely to be physically realistic.
Figure 10: The orbital period distributions of the 2017ApJS..230...15M, 1991A&A...248..485D, and 1992ApJ...396..178F models, respectively.
@@
Figure 11: Top: The PPDs derived from the 1991A&A...248..485D (blue) and 1992ApJ...396..178F (red) models, where the only difference is the shift in the mean value of the log-normal period distribution from 4.8 (1991A&A...248..485D) to 3.5 (1992ApJ...396..178F). This change causes the binary fraction to decrease significantly, from 0.61 to 0.38, as expected. Bottom: The PPDs for the metal-rich (orange/cyan) and metal-poor (green/purple) subsamples under the 1991A&A...248..485D and 1992ApJ...396..178F models, respectively. When a lower mean period is applied, the binary fractions for both subsamples are consistently shifted to lower values.
4.4 Is one year of time baseline enough?
@@
We first assess the fraction of binary systems that can be detected with the one-year observational baseline. To achieve this, we generate a mock sample of binaries by randomly selecting 1,000 samples from our simulations (see Section 3.2), which contains about 670,000 stars with over 2 million exposures in total. The large sample size ensures that statistical fluctuations are small enough. In this sample, the binary fraction is set to 1. Besides, we do not include any Gaussian noise, which is described in step 4 of Section 3.2, to isolate the intrinsic LOSV variations purely induced by the orbital motion of the binaries.
We use the standard deviation of the LOSV, $\sigma_{v_{\rm{los}}}$, to quantify velocity variations in this test. The results from the mock binary sample are shown in Figure 12, where the distribution of $\sigma_{v_{\rm{los}}}$ is represented by the blue histogram. As illustrated in the figure, approximately 12% of the simulated binaries exhibit LOSV variations exceeding 1.2  km s-1, as indicated by the black vertical dashed line. This threshold represents the systematic error floor of LOSV measurements of the member stars in UMi from the rvs pipeline, as mentioned in Section 2. Due to the limited one-year observational baseline, the LOSV variations of long-period binaries remain small. Examining the three additional histograms, which divide the sample based on orbital periods, we observe a clear trend. The majority ($99.6\%$) of long-period binaries ($P>10$ years, red histogram) exhibit velocity variations of $\sigma_{v_{\rm{los}}}<1.2$ km $s^{-1}$, making them difficult to detect. In contrast, short-period binaries ($P<1$ year, orange histogram) show significantly larger velocity variations, while the medium-period binaries ($1<P<10$ years, green histogram) fall between these two groups, displaying moderate velocity variations. Among the entire binary sample, about 12% of stars with LOSV variations surpassing the 1.2 km s-1 error floor provide meaningful constraints on the binary fraction. The remaining stars, with LOSV variations below this threshold, are dominated by uncertainties and contribute negligibly to our analysis.
@@
In our test associated with Figure 13, however, we adopted exactly the same binary orbital parameter distribution model used for generating the mock data to recover the binary fraction. So the model extrapolation is perfect. If, in reality, the underlying binary orbital parameter distributions differ from what we assume of using the 1991A&A...248..485D or 2017ApJS..230...15M models, the extrapolation may not work as perfectly as what we have demonstrated with Figure 13 to recover the actual binary fractions. Nevertheless, we can at least conclude that for short-period binaries that can be detected within a one-year time baseline, the metal-rich population has slightly higher binary fractions than the metal-poor population. Future observations, including more dedicated tiles of DESI MWS, can further increase the time baselines and will be crucial for obtaining more comprehensive and precise constraints on the binary fractions for longer period binaries in UMi, in other dwarf galaxies, in globular clusters, and in the field of our MW.
@@ 5 Discussions
Our analysis reveals that the metal-poor stellar population in UMi exhibits a slightly lower binary fraction compared to the metal-rich population. Given the results of the selection effect analysis in Section 4.2 and the discussion in Section 4.3, we can rule out observational biases as the primary drivers of this detection. To better understand these results, it is useful to compare them with previous studies that have explored binary fraction trends across different stellar populations.
@@
Several studies have investigated how binary fractions vary with stellar properties such as mass, metallicity, and age. For instance, 2010ApJS..190....1R conducts a comprehensive investigation of companions to solar-type stars up to 25 pc, showing that more massive, bluer stars are more likely to host binary companions than redder, lower-mass stars. Additionally, they find that among redder stars, metal-poor systems exhibit a higher probability of hosting companions. Similarly, 2018RAA....18...52T examines the relationship between binary fraction and $\alpha$-element abundances. They find that older stars with higher $\alpha/\mathrm{Fe}]$ tend to have a higher binary fraction. 2018ApJ...854..147B investigated the multiplicity of field stars. They also reveal a metallicity dependency, with metal-poor stars ($\mathrm{Fe/H}]\lesssim-0.5$) exhibiting a multiplicity that is 2–3 times higher than that of metal-rich stars ($\mathrm{Fe/H}]\gtrsim 0.0$). Besides, the study by 2019ApJ...875...61M finds that the close binary fraction ($P<10^{4}$ days; $a<10$ au) among solar-type stars decreases as the metallicity increases. Another work by 2019MNRAS.484.2341B reports a similar trend from a few radiation hydrodynamical simulations, that the multiplicity slightly increases as metallicity decreases, which is due to the greater cooling rates at high gas densities due to the lower opacities at low metallicities increasing the gas fragmentation on small scales. These studies all suggest that stars formed earlier, which are typically more metal-poor and $\alpha$-enhanced, have higher binary fractions.
@@
It seems our analysis reveals a slightly opposite trend from those previous studies: stars in UMi with lower [Fe/H] and thus older, have smaller binary fractions. However, direct comparisons between our results and previous studies are not straightforward. The metal-poor populations in these previous works are not sufficiently metal-poor to match the conditions of UMi. As shown in the lower left panel of Figure 3, even the most metal-rich member star in our UMi sample has $\mathrm{[Fe/H]}\simeq-1$, reflecting the extremely low metallicity environment in UMi. Our analysis focuses on a regime that is significantly more metal-poor than those in earlier studies. As a comparison, the most metal-poor stars in 2018RAA....18...52T have $\mathrm{[Fe/H]}>-1.0$, a value that would be considered part of the metal-rich population in our analysis. Although the stars analyzed in 2018ApJ...854..147B span a broad metallicity range down to $\mathrm{Fe/H}]\approx-2.5$, the majority is from -1 to 0.5, and their metal-rich sample is defined as $\mathrm{Fe/H}]\gtrsim 0.0$. Similarly, Figure 15 of 2019MNRAS.484.2341B shows that their sample only covers metallicities greater than 0.01 $\rm{Z}_{\odot}$. In the case of 2010ApJS..190....1R, they only include stars with $\mathrm{[Fe/H]}>-0.9$, again significantly more metal-rich than our case. Furthermore, most studies primarily focus on solar-type stars around our neighborhood, where the average metallicity is already enhanced compared to that of dwarf galaxies. Given these differences, we cannot directly apply their conclusions to our study.
@@
A crucial factor to consider might be the significantly older age of the stellar population in our sample than those studied in previous works. Dynamical evolution likely plays a role in shaping the observed binary fraction. Older, metal-poor stars have undergone a longer time for dynamical interactions to happen, increasing the risk of disruption due to encounters with flyby stars or other tidal interactions. As suggested by previous works (2010arXiv1005.5388P; 2016MNRAS.461L..72P; 2022ApJ...929...77K), wide stellar binaries in dwarf spheroidal galaxies can be disrupted by repeated encounters with dark substructures orbiting within their dark matter halos. The much more extended evolutionary history for older stars formed in dwarf galaxies may explain the lower binary fraction observed in metal-poor populations. For the other previous studies mentioned above, their sample of stars may not have undergone evolutionary histories that are as old as member stars in old dwarf galaxies, so they may not have enough time to accumulate enough dynamical encounters to disrupt binary systems.
@@
However, as we have mentioned in Section 4.4 above, our one-year time baseline makes it hard to detect the LOSV variability of binaries with orbital periods longer than 10 years, for which the best-constrained binary fractions rely on model extrapolations. For binaries with orbital periods of 1 to 10 years that can be more safely detected by our time window, the typical binary separation is at most a few AU, assuming a circular orbit and a mass ratio of 1. However, given the typical density of stars in MW dwarf galaxies, we estimate the typical mean free path of close encounters to be 1000 AU in UMi, assuming the same stellar number density as in Section 3.1.3. This suggests that the separations of short-period binaries that can be detected by our observations are too small to be sensitive to close encounters. Thus, the explanation of long-time dynamical interactions to disrupt binaries in the metal-poor population of dwarf galaxies may only be valid if our model extrapolations hold for longer-period binaries.
Another explanation associated with the disruption of short-period binaries may be related to the orbital decay and eventual merger driven by dynamical friction in their early evolutionary phase. When young binaries form within dense molecular clouds, they experience drag from the surrounding gas, which carries away angular momentum through acoustic disturbances in the form of spiral waves (e.g. 2010MNRAS.402.1758S).
@@
In the end, we mention recent studies by 2025AJ....170..171L and 2025A&A...702A..11G, which provide independent evidence that is more consistent with our findings. 2025AJ....170..171L identified nearly equal-mass binaries through the difference between the geometric absolute magnitudes from Gaia parallax and the absolute magnitudes inferred from spectrophotometric distances measured by a data-driven approach. Their results indicate a positive correlation between binary fraction and metallicity, consistent with our findings in UMi. Moreover, a recent analysis of BHB stars by 2025A&A...702A..11G found that halo-like, metal-poor populations exhibit lower intrinsic binary fractions than their more metal-rich, disk-like counterparts. This contrast was interpreted as evidence that halo BHBs likely formed predominantly through single-star evolutionary channels, while disk BHBs experienced more binary-driven evolution. Their finding aligns broadly with our result.
@@
6 Conclusion
In this paper, we conduct a detailed investigation of the binary fraction in the Ursa Minor dwarf galaxy (UMi) using spectroscopic data from the DESI Milky Way Survey. Our dataset consists of multiple exposures taken over approximately one year, providing a large sample of LOSV measurements that enable us to track LOSV variability as a proxy for binary abundance.
Adopting the orbital parameters from 1991A&A...248..485D (hereafter 1991A&A...248..485D) and 2017ApJS..230...15M (hereafter 2017ApJS..230...15M) models, we derive overall binary fractions for UMi of $0.61^{+0.16}_{-0.20}$ and $0.69^{+0.19}_{-0.17}$, respectively. These estimates are consistent with previous studies, demonstrating that high binary fractions are present in UMi.
After dividing member stars in UMi into two approximately equal-sized subsamples by metallicity, we find a discrepancy between the inferred binary fractions of the metal-rich and metal-poor populations, with metal-rich stars exhibiting slightly higher binary fractions. In the 1991A&A...248..485D model, metal-rich sample reaches $0.72^{+0.23}_{-0.17}$, while metal-poor sample drops to $0.46^{+0.24}_{-0.19}$. Similarly, with the 2017ApJS..230...15M model, these values are $0.86^{+0.14}_{-0.24}$ and $0.48^{+0.26}_{-0.19}$ for the metal-rich and metal-poor subsamples, respectively.

### v2

@@ IV.1 Binary Fraction in UMi
Figure 7: We present the PPDs for the metal-rich (orange) and metal-poor (green) populations, as indicated in the legend. The top and bottom panels are based on the DM91 and MS17 models, respectively. The blue curves show the results for the full sample (as in Figure 6).
IV.2 Selection Effects
In Section IV.1, we report a marginal indication that the metal-rich stellar population of UMi may exhibit a higher binary fraction than the metal-poor population. In this section, we examine whether this metallicity dependency could arise from sample selection effects rather than reflecting an intrinsic difference. Specifically, we investigate the impact of non-uniform observing strategies across UMi, including variations in the number of epochs and the time baselines or separations between them. For example, we know that the spatial distributions of metal-rich and metal-poor stars differ, with more metal-rich stars in central regions. The tiling of UMi, however, is spatially varying (see Figure 1), which may result in different number of epochs and time baselines for different stellar populations and may affect the comparison between different populations.
@@ IV.2 Selection Effects
Figure 8: Resampling of the metal-rich and metal-poor populations to match their distributions in number of epochs (left column), distance to the dwarf center in projection (semi-major axis of elliptical isophotes in number density, middle column), and LOSV uncertainty (right column), respectively. Top: Original distributions for the metal-rich (orange) and metal-poor (green) samples, along with the matched distribution after resampling (dashed red). Middle: The PPDs of the binary fraction for the DM91 model following resampling in both populations. Bottom: The corresponding results for the MS17 model. We perform the resampling procedure 100 times. The translucent histograms show the distribution of results from individual realizations, while the bold lines represent the combined PPDs from all trials.
@@
The middle (DM91 model) and bottom (MS17 model) rows of Figure 8 show the PPDs of the binary fractions for the resampled metal-rich (orange) and metal-poor (green) populations. Note that the solid curve is the overall PPD from the 100 resampled realizations mentioned above, with the results for the 100 realizations shown in the background by curves with thin, lighter colored lines. Although the resampling slightly alters the measured binary fractions and increases the uncertainties due to the reduced sample size, the difference between the binary fractions of the metal-rich and metal-poor populations remains.
@@
Figure 9: Similar to Figure 8, we resample both the metal-rich and metal-poor subsamples to ensure they share the same distribution of time separations across all observation pairs. Top: The distribution of the time separation in units of day for individual pairs of observations. Middle: The PPDs derived from the DM91 model after resampling. The orange histogram is the result of the metal-rich population, while the green one is the result of the metal-poor population. Bottom: The corresponding results from the MS17 model. Both models show consistent results, with the discrepancy between the two subsamples persisting despite the resampling.
@@
IV.3 Systematics Associated with the Period Distribution
@@
In addition to observational factors, the inferred binary fraction is highly sensitive to the choice of model parameter distributions. As mentioned in Section III.1, the orbital period distribution plays a major role in constraining the binary fraction, with a lower mean value leading to a substantial decrease in the inferred binary fraction.
@@
In Section IV.1, the binary fraction derived using the MS17 model is slightly higher than that obtained with the DM91 model. The key distinction between these models lies in their treatment of orbital parameter distributions. While the DM91 model assumes independent orbital parameter distributions, the MS17 model provides joint distributions that account for correlations between different orbital parameters. We present the period distributions in Figure 10 for different models. Compared to the DM91 model, represented by the orange line, the MS17 model, shown by the blue line, predicts a bit smaller fraction of short-period binaries. This is the major reason that leads to an overall slight increase in the inferred binary fraction by the MS17 model, as demonstrated in our results above.
@@
This sensitivity is also demonstrated in S18, which infers that the binary fraction decreases as the mean period value of the model, $\mu_{\log P}$, increases. S18 uses three different models, including Fischer and Marcy (1992) (hereafter FM92), DM91 and Marks and Kroupa (2011), which differ only in the assumed mean period values of $\mu_{\log P}$ = 3.5, 4.8, and 5.8, respectively. The corresponding binary fractions derived from these models by S18 are 0.52, 0.78, and 0.96, illustrating the strong dependency of the inferred binary fraction on the choice of period distribution.
@@
To further investigate the impact of the orbital period distribution, we repeat our constraints adopting the mean period value from the FM92 model [^5] We have also explored the Marks and Kroupa (2011) period distribution model, which leads to a binary fraction very close to one and might not be realistic, so we do not show the results based on this model., which is lower than the value of DM91, while keeping all other orbital parameter distributions identical to those in DM91. For clarity, we refer to this modified version as the “FM92” model for our discussions afterward. The green line in Figure 10 represents the orbital period distribution for the FM92 model, which shows a higher fraction of shorter-period binaries compared to the DM91 model. As shown in the upper panel of Figure 11, the binary fraction derived from the FM92 model decreases by approximately 0.2 for the full sample compared with that from DM91. This difference arises primarily due to the shift towards shorter orbital periods, and thus a lower binary fraction is preferred for the same LOSV variability.
@@
The bottom panel of Figure 11 shows that the binary fractions for both the metal-rich and metal-poor subsamples decrease by approximately the same amount when applying the FM92 model with a lower mean period. Despite the decrease in the inferred binary fraction, the discrepancy between the metal-rich and metal-poor populations persists (see the cyan and purple curves). Here we emphasize that we are applying the same orbital period distribution model to both metal-rich and metal-poor populations. If the metal-rich and metal-poor populations follow different period distributions, such as a shorter mean period for the metal-rich population and a longer mean period for the metal-poor population, the observed discrepancy in binary fractions could be reduced or even eliminated. This effect can be seen by comparing the green curve of the metal-poor population and the cyan curve for the metal-rich population in the bottom panel of Figure 11. However, explaining the discrepancy in this way would require the metal-rich and metal-poor populations to have fundamentally different orbital period distributions, for which there is no strong theoretical support and which is therefore unlikely to be physically realistic.
Figure 10: The orbital period distributions of the MS17, DM91, and FM92 models, respectively.
@@
Figure 11: Top: The PPDs derived from the DM91 (blue) and FM92 (red) models, where the only difference is the shift in the mean value of the log-normal period distribution from 4.8 (DM91) to 3.5 (FM92). This change causes the binary fraction to decrease significantly, from 0.61 to 0.38, as expected. Bottom: The PPDs for the metal-rich (orange/cyan) and metal-poor (green/purple) subsamples under the DM91 and FM92 models, respectively. When a lower mean period is applied, the binary fractions for both subsamples are consistently shifted to lower values.
IV.4 Is one year of time baseline enough?
@@
We first assess the fraction of binary systems that can be detected with the one-year observational baseline. To achieve this, we generate a mock sample of binaries by randomly selecting 1,000 samples from our simulations (see Section III.2), which contains about 670,000 stars with over 2 million exposures in total. The large sample size ensures that statistical fluctuations are small enough. In this sample, the binary fraction is set to 1. Besides, we do not include any Gaussian noise, which is described in step 4 of Section III.2, to isolate the intrinsic LOSV variations purely induced by the orbital motion of the binaries.
We use the standard deviation of the LOSV, $\sigma_{v_{\rm{los}}}$, to quantify velocity variations in this test. The results from the mock binary sample are shown in Figure 12, where the distribution of $\sigma_{v_{\rm{los}}}$ is represented by the blue histogram. As illustrated in the figure, approximately 12% of the simulated binaries exhibit LOSV variations exceeding 1.2  km s-1, as indicated by the black vertical dashed line. This threshold represents the systematic error floor of LOSV measurements of the member stars in UMi from the rvs pipeline, as mentioned in Section II. Due to the limited one-year observational baseline, the LOSV variations of long-period binaries remain small. Examining the three additional histograms, which divide the sample based on orbital periods, we observe a clear trend. The majority ($99.6\%$) of long-period binaries ($P>10$ years, red histogram) exhibit velocity variations of $\sigma_{v_{\rm{los}}}<1.2$ km $s^{-1}$, making them difficult to detect. In contrast, short-period binaries ($P<1$ year, orange histogram) show significantly larger velocity variations, while the intermediate-period binaries ($1<P<10$ years, green histogram) fall between these two groups, displaying moderate velocity variations. Among the entire binary sample, about 12% of stars with LOSV variations surpassing the 1.2 km s-1 error floor provide meaningful constraints on the binary fraction. The remaining stars, with LOSV variations below this threshold, are dominated by uncertainties and contribute negligibly to our analysis.
@@
In our test associated with Figure 13, however, we adopted exactly the same binary orbital parameter distribution model used for generating the mock data to recover the binary fraction. So the model extrapolation is perfect. If, in reality, the underlying binary orbital parameter distributions differ from what we assume of using the DM91 or MS17 models, the extrapolation may not work as perfectly as what we have demonstrated with Figure 13 to recover the actual binary fractions. Nevertheless, we can at least conclude that for short-period binaries that can be detected within a one-year time baseline, the metal-rich population has slightly higher binary fractions than the metal-poor population. Future observations, including more dedicated tiles of DESI MWS, can further increase the time baselines and will be crucial for obtaining more comprehensive and precise constraints on the binary fractions for longer period binaries in UMi, in other dwarf galaxies, in globular clusters, and in the field of our MW.
In addition, these simulations allow us to directly assess the impact of binaries on the measured velocity dispersion. By analyzing the mock binary sample described above and assuming a binary fraction of 1, we derive the velocity dispersion induced solely by binary orbital motions. As we mentioned above, long-period binaries ($P>10$ years) produce velocity variations that are indistinguishable from measurement uncertainties and therefore contribute negligibly to the likelihood. Consistently, their impact on the inferred velocity dispersion is also minimal, at the level of less than 1%, which does not affect our subsequent analysis. Consequently, even in the extreme case of a high intrinsic binary fraction (assumed to be 1), long-period binaries do not inflate the total velocity dispersion; instead, the dominant contribution arises from short- and intermediate-period systems, which could increase the dispersion by approximately 17% if all binaries are assumed to have periods shorter than 10 years. When binaries of all periods are included, the overall inflation is about 4%.
@@ V Discussions
Our analysis reveals that the metal-poor stellar population in UMi exhibits a slightly lower binary fraction compared to the metal-rich population. Given the results of the selection effect analysis in Section IV.2 and the discussion in Section IV.3, we can rule out observational biases as the primary drivers of this detection. To better understand these results, it is useful to compare them with previous studies that have explored binary fraction trends across different stellar populations.
@@
Several studies have investigated how binary fractions vary with stellar properties such as mass, metallicity, and age. For instance, Raghavan et al. (2010) conducts a comprehensive investigation of companions to solar-type stars up to 25 pc, showing that more massive, bluer stars are more likely to host binary companions than redder, lower-mass stars. Additionally, they find that among redder stars, metal-poor systems exhibit a higher probability of hosting companions. Similarly, Tian et al. (2018) examines the relationship between binary fraction and $\alpha$-element abundances. They find that older stars with higher $\alpha/\mathrm{Fe}]$ tend to have a higher binary fraction. Badenes et al. (2018) investigated the multiplicity of field stars. They also reveal a metallicity dependency, with metal-poor stars ($\mathrm{Fe/H}]\lesssim-0.5$) exhibiting a multiplicity that is 2–3 times higher than that of metal-rich stars ($\mathrm{Fe/H}]\gtrsim 0.0$). Besides, the study by Moe et al. (2019) finds that the close binary fraction ($P<10^{4}$ days; $a<10$ au) among solar-type stars decreases as the metallicity increases. Another work by Bate (2019) reports a similar trend from a few radiation hydrodynamical simulations, that the multiplicity slightly increases as metallicity decreases, which is due to the greater cooling rates at high gas densities due to the lower opacities at low metallicities increasing the gas fragmentation on small scales. These studies all suggest that stars formed earlier, which are typically more metal-poor and $\alpha$-enhanced, have higher binary fractions.
@@
It seems our analysis reveals a slightly opposite trend from those previous studies: stars in UMi with lower [Fe/H] and thus older, have smaller binary fractions. However, direct comparisons between our results and previous studies are not straightforward. The metal-poor populations in these previous works are not sufficiently metal-poor to match the conditions of UMi. As shown in the lower left panel of Figure 3, even the most metal-rich member star in our UMi sample has $\mathrm{[Fe/H]}\simeq-1$, reflecting the extremely low metallicity environment in UMi. Our analysis focuses on a regime that is significantly more metal-poor than those in earlier studies. As a comparison, the most metal-poor stars in Tian et al. (2018) have $\mathrm{[Fe/H]}>-1.0$, a value that would be considered part of the metal-rich population in our analysis. Although the stars analyzed in Badenes et al. (2018) span a broad metallicity range down to $\mathrm{Fe/H}]\approx-2.5$, the majority is from -1 to 0.5, and their metal-rich sample is defined as $\mathrm{Fe/H}]\gtrsim 0.0$. Similarly, Figure 15 of Bate (2019) shows that their sample only covers metallicities greater than 0.01 $\rm{Z}_{\odot}$. In the case of Raghavan et al. (2010), they only include stars with $\mathrm{[Fe/H]}>-0.9$, again significantly more metal-rich than our case. Furthermore, most studies primarily focus on solar-type stars around our neighborhood, where the average metallicity is already enhanced compared to that of dwarf galaxies. Given these differences, we cannot directly apply their conclusions to our study.
@@
A crucial factor to consider might be the significantly older age of the stellar population in our sample than those studied in previous works. Dynamical evolution likely plays a role in shaping the observed binary fraction. Older, metal-poor stars have undergone a longer time for dynamical interactions to happen, increasing the risk of disruption due to encounters with flyby stars or other tidal interactions. As suggested by previous works (Peñarrubia et al., 2010, 2016; Kervick et al., 2022), wide stellar binaries in dwarf spheroidal galaxies can be disrupted by repeated encounters with dark substructures orbiting within their dark matter halos. The much more extended evolutionary history for older stars formed in dwarf galaxies may explain the lower binary fraction observed in metal-poor populations. For the other previous studies mentioned above, their sample of stars may not have undergone evolutionary histories that are as old as member stars in old dwarf galaxies, so they may not have enough time to accumulate enough dynamical encounters to disrupt binary systems.
@@
However, as we have mentioned in Section IV.4 above, our one-year time baseline makes it hard to detect the LOSV variability of binaries with orbital periods longer than 10 years, for which the best-constrained binary fractions rely on model extrapolations. For binaries with orbital periods of 1 to 10 years that can be more safely detected by our time window, the typical binary separation is at most a few AU, assuming a circular orbit and a mass ratio of 1. However, given the typical density of stars in MW dwarf galaxies, we estimate the typical mean free path of close encounters to be 1000 AU in UMi, assuming the same stellar number density as in Section III.1.3. This suggests that the separations of short-period binaries that can be detected by our observations are too small to be sensitive to close encounters. Thus, the explanation of long-time dynamical interactions to disrupt binaries in the metal-poor population of dwarf galaxies may only be valid if our model extrapolations hold for longer-period binaries.
Another explanation associated with the disruption of short-period binaries may be related to the orbital decay and eventual merger driven by dynamical friction in their early evolutionary phase. When young binaries form within dense molecular clouds, they experience drag from the surrounding gas, which carries away angular momentum through acoustic disturbances in the form of spiral waves (e.g. Stahler, 2010).
@@
In the end, we mention recent studies by Li et al. (2025) and Guo et al. (2025), which provide independent evidence that is more consistent with our findings. Li et al. (2025) identified nearly equal-mass binaries through the difference between the geometric absolute magnitudes from Gaia parallax and the absolute magnitudes inferred from spectrophotometric distances measured by a data-driven approach. Their results indicate a positive correlation between binary fraction and metallicity, consistent with our findings in UMi. Moreover, a recent analysis of BHB stars by Guo et al. (2025) found that halo-like, metal-poor populations exhibit lower intrinsic binary fractions than their more metal-rich, disk-like counterparts. This contrast was interpreted as evidence that halo BHBs likely formed predominantly through single-star evolutionary channels, while disk BHBs experienced more binary-driven evolution. Their finding aligns broadly with our result.
@@
VI Conclusion
In this paper, we conduct a detailed investigation of the binary fraction in the Ursa Minor dwarf galaxy using spectroscopic data from the DESI Milky Way Survey. Our dataset consists of multiple exposures taken over approximately one year, providing a large sample of LOSV measurements that enable us to track LOSV variability as a proxy for binary abundance.
Adopting the orbital parameters from Duquennoy and Mayor (1991) (hereafter DM91) and Moe and Di Stefano (2017) (hereafter MS17) models, we derive overall binary fractions for UMi of $0.61^{+0.16}_{-0.20}$ and $0.69^{+0.19}_{-0.17}$, respectively. These estimates are consistent with previous studies, demonstrating that high binary fractions are present in UMi.
After dividing member stars in UMi into two approximately equal-sized subsamples by metallicity, we find a discrepancy between the inferred binary fractions of the metal-rich and metal-poor populations, with metal-rich stars exhibiting slightly higher binary fractions. In the DM91 model, metal-rich sample reaches $0.72^{+0.23}_{-0.17}$, while metal-poor sample drops to $0.46^{+0.24}_{-0.19}$. Similarly, with the MS17 model, these values are $0.86^{+0.14}_{-0.24}$ and $0.48^{+0.26}_{-0.19}$ for the metal-rich and metal-poor subsamples, respectively.